Two consecutive odd numbers have a sum of 44. Find the value of the product of both of these odd numbers?
Answer by macston(5194)

First number is X, consecutive odd number is X+2
X+X+2=44
2X+2=44
2X=42
X=21 First number is 21
X+2=23 Consecutive odd number is 23
Product=21(23)=483
The product of two consecutive odd numbers with a sum of 44 is 483